OpenStudy (anonymous):

use algebraic equation to find the measures of the 2 angles described,,, begin by letting x represent the degree measure of the angles supplement. the measure of the angle is 59 times greater than its supplement what is the measure of the supplement what is the measure of the other angel

OpenStudy (anonymous):

x +59 x=180

OpenStudy (anonymous):

whats the measure of the supplement what is the measure of the angle

OpenStudy (anonymous):

supplement is 3 nd angle is 59*3

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

welcum

OpenStudy (anonymous):

may i ask u another question of should i just put it out there for all

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

find the coordinates of the other endpoint of the segment, given its mid point and one endpoint. midpoint (2,-2) end point(-4,-11) what is the other endpoint

OpenStudy (anonymous):

mid point formula is x= x1+x2/2 nd y=y1 +y2/2... u got x nd x1.. find x2... similarly find y2

OpenStudy (anonymous):

is it (-1,-7.5)

OpenStudy (anonymous):

no its 8 nd 7

OpenStudy (anonymous):

how is it 8 and 7